result in simplest form

3 1/3

divide


2 2/3

Answer by fractalier(6550)

3 1/3 divided by 2 2/3
Let us change both to improper fractions first...so we get
10/3 divided by 8/3
Then flip (or invert) the second number and change it to a multiplication problem...giving us
10/3 * 3/8 =
now cancel threes and get
10/8 = 5/4 or 1 1/4